OpenStudy (anonymous):

I need help to work this problem: 1/10-1/5?

OpenStudy (amistre64):

it is the same as: 1/10 - 2/10 = -1/10

OpenStudy (anonymous):

Thanks. How about this:-4y+9=-5y+7?

OpenStudy (amistre64):

combine like terms: -4y + 5y = 7 - 9 y = -2